How many liters each of 25% and 60% alcohol solutions must be mixed together to make 40 liters of a 53% alcohol solution?

Answer by TimothyLamb(4379)

w = weak solution volume
s = strong solution volume
---
alcohol in the mixed solution:
0.25w + 0.60s = 0.53*40
---
total volume in the mixed solution:
w + s = 40
w = 40 - s
---
0.25w + 0.60s = 0.53*40
0.25(40 - s) + 0.60s = 21.2
10 - 0.25s + 0.60s = 21.2
0.35s = 11.2
---
Answer:
s = 32 liters
w = 8 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php